OpenStudy (anonymous):

The crest of the second hill is circular, with a radius of r = 35.5 m. Neglect friction and air resistance. What must be the height h of the first hill so that the skier just loses contact with the snow at the crest of the second hill?

OpenStudy (anonymous):

When the skier loses contact on the second hill,the N(normal reaction due to hill) vanishes.Consider the forces acting on skier at the top of the hill. 1]Mg acts downwards 2]N upwards Therefore, \[Mg = m \frac{V^2}{r}\] There is no N in the above equation because it vanishes when the skier leaves the contact Therefore: \[Vf = \sqrt{g*r}\] Now apply Energy conservation at 2 pts at the top of two hills: \[mgH = mg(35.5) + \frac{1}{2}m(Vf)^2\] makes sense?

OpenStudy (anonymous):

Solving for H at the end

OpenStudy (anonymous):

u are the real man :)

OpenStudy (anonymous):

thanks hehe

OpenStudy (anonymous):

do they ask what the mass is ?

OpenStudy (anonymous):

cause it's not needed, the mass cancels out oin the equation

OpenStudy (anonymous):

ok